OpenStudy (anonymous):

the second term in a geometric sequence is 50. the fourth term in the same sequence is 112.5. what is the common ratio in this sequence?

OpenStudy (anonymous):

a, 50, b, 112.5 50/a = b/50 = 112.5/b use those equations above to solve for a and b, then find the common ratio

OpenStudy (anonymous):

I cant think of what will replace a and b

OpenStudy (whpalmer4):

actually it's even simpler: \[\frac{b}{50} = \frac{112.5}{b}\]

OpenStudy (tkhunny):

You do not need to replace "a". It is not needed at all for this problem.

OpenStudy (whpalmer4):

then your common ratio is just \[\frac{112.5}{b}\] or \[\frac{b}{50}\]

OpenStudy (anonymous):

is there a process to get b?

OpenStudy (whpalmer4):

or better yet, \[50*a*a=112.5\]where \(a\) is your common ratio

OpenStudy (whpalmer4):

cross multiply and solve for \(b\) if you're going with the \[\frac{b}{50}=\frac{112.5}{b}\]formulation. \[b*b = 50*112.5\]

OpenStudy (anonymous):

i got 75 and the ratio came to be 1.5... i guess that's right

OpenStudy (tkhunny):

Prove IT!!! a = 1st term r = common ratio ar = 2nd term = 50 ar^2 = 3rd term ar^3 = 4th term = 112.5 There you go! Now, solve for r. ar = 2nd term = 50 ==> a = 50/r ar^2 = 3rd term ar^3 = 4th term = 112.5 ==> a = 112.5/(r^3) 50/r = 112.5/(r^3)